Daily Report for IODP Expedition 311, September 20, 2005


LOCATION: Site U1325 (CAS-02C)


SCIENCE UPDATE: We arrived on site, switched to DP mode, settling on location at 07:25. While deploying the LWD tool string, we determined that the ProVision NMR tool was not communicating to the surface.  However, the ProVision began to function after spudding the hole. A “tool-box” meeting was held to review LWD protocols immediately prior to initiating drilling. Hole U1325A was spudded at 23:10.


Planning for coring progressed today with a general science meeting to discuss core flow and sampling procedures. Subgroup meetings to discuss lab procedures and sample requests continued. An integrated gas hydrate sampling plan was developed and circulated to the scientific and technical staff to address sampling protocols on the catwalk including safety issues for H2S and core handling hazards.